She needs to prepare 36oz of a 12% solution. Find the amount of 18% solution and the amount of 9% solution she should mix to get this solution.
Answer by ankor@dixie-net.com(22740)

She needs to prepare 36oz of a 12% solution.
Find the amount of 18% solution and the amount of 9% solution she should mix to get this solution.
:
Let x = amt of 18% solution
the result is to be 36oz, therefore
(36-x) = amt of 9% solution
:
A mixture equation in decimal form
.18x + .09(36-x) = .12(36)
.18x + 3.24 - .09x = 4.32
.18x - .09x = 4.32 - 3.24
.09x = 1.08
x = 1.08/.09
x = 12 oz of 18%, then obviously it's 24 oz of 9%
;
:
you can check this on your calc
.18(12) + .09(24) = .12(36)